Two jet leave Atlanta in opposite directions at th same time. The speed of the jet headed east is two times of the jet headed west. In three ours they are 1089 miles apart. Find the speed of each jet.
Answer by rfer(16322)

1089/3=363
2*363=726/3=242 mph
363/3=121 mph